Ziggy's famous yogurt b lends regular yogurt that is 3% fat with its no-fat yogurt to obtain low fat yogurt that is 1% fat. How many pounds of regular yogurt and how many pounds of no fat yougurt should be mixed to obtain 60 pounds of low fat yogurt?
Answer by ptaylor(2198)

Let x=number of pounds 3% fat yogurt needed
Then 60-x=number of pounds of no-fat yogurt needed
Now we know that the amount of pure fat in the 3% yogurt (0.03x) plus the amount of pure fat in the no-fat yogurt(0) must equal to the amount of pure fat in the final yogurt mixture ((0.01)(60)). So our equation to solve:
0.03x=0.01*60 or
0.03x=0.6 divide both sides by 0.03
x=20 lb-------------------------------number of pounds of 3% fat yogurt needed
60-x=60-20=40 lb -----------------number of pounds of no-fat yogurt needed
CK
0.03*20=0.01*60
0.6=0.6
Hope this helps----ptaylor